OpenStudy (anonymous):

MEDAL AND FAN! Solve for x: 4|2x-2| + 10 =34 A) x=-4, x=2 B) x=4, x=-4 C) x=2, x=-2 D) x=4, x=-2

OpenStudy (anonymous):

@misty1212 can you help me with this one too?

OpenStudy (anonymous):

@SolomonZelman can you help me?

OpenStudy (misty1212):

\[[2|2x-2|+10=34\]subtract 10 from both sides

OpenStudy (misty1212):

\[4|2x-2|=24\]

OpenStudy (misty1212):

\[8|x-1|=24\] is good too then divide by \(8\) what do you get?

OpenStudy (anonymous):

The absolute value of x-1 = 3

OpenStudy (misty1212):

good then you have to solve \px-1=3\] and also \[x-1=-3\] both are by adding one to both sides

OpenStudy (misty1212):

\[x-1=3\] and \[x-1=-3\]

OpenStudy (misty1212):

add one to each side for two answers

OpenStudy (anonymous):

so x=4 and x=-2 =3

OpenStudy (anonymous):

^_^ Thankies so much @misty1212

OpenStudy (misty1212):

so yes \(x=4\) or \(x=-2\)

OpenStudy (misty1212):

welkies again thanks for the medal

OpenStudy (triciaal):

another approach first -10 from each side then divide each side by 4 now you have 2 equations because you have an absolute on the left , you will take the positive = right hand side and solve on the left now take the negative = right hand side and solve

OpenStudy (misty1212):

;)